Fermi National Laboratory


Accelerator Update

Friday August 15
The day shift began with Operations monitoring store 2910, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 9:42 AM, a Linac RF station (LRF2) tripped off and wouldn't reset. Linac technicians began troubleshooting the problem.

Operators made Booster ready for an access at 10:26 AM to investigate possible vacuum problems. The access began at 10:44 AM.

Operators also made the Main Injector (MI) and Pbar ready for access at 10:12 AM. MI techs were out of access by 2 PM and Pbar out by 3 PM.

MiniBooNE went in on access at 10:48 AM. They were out by 2:30 PM.

Experts accessed the Linac at 11:50 AM. They were out thirty minutes later.

Vacuum techs reported finding and repairing the Booster vacuum leak at 12:32 PM.

Operations established beam to the Linac dump at 3:38 PM.

The Booster and all other areas were out of access by 4 PM and being turned on.

Operations established beam to Booster at 4:31 PM, and to MI, MiniBooNE, and Pbar by 5 PM.

LRF2 tripped off for the first of three times over twenty minutes at 5 PM.

Another Linac RF station (KRF6) tripped off at 5:04 PM. An expert investigated. He had the station working by 5:20 PM.

A Tevatron (TeV) sector B3 vacuum pump tripped off at 7:13 PM and would not reset.

Saturday August 16
The midnight shift began with Operations monitoring store 2910, with Pbar stacking, and with MiniBooNE taking beam.

At 12:42 AM, LRF2 tripped off. Operators responded, but could not get the station back on. This held off beam to all other accelerators. Linac techs arrived about 2:30 AM. They found a bad relay and had it replaced and the station back on by 3:53 AM.

LRF2 tripped off several more times between 4 and 5 AM.

Operations terminated store 2910 at 6:19 AM. The TeV was in shot setup by 6:56 AM.

Operations established store 2912 at 9:29 AM with an initial luminosity of 3.8E31. Pbar resumed stacking.

A Recycler expert began taking shots of protons, for study purposes, at 11:08 AM.

At 2 PM, the TeV suffered a sector F4 quench, which aborted the store. Operations put the TeV into recovery. After much investigation experts believed that a controls problem may have caused the abort, but this was not certain. Experts did note a vertical distortion in the beam orbit. Operations called in controls experts.

Operations put the TeV into a dry squeeze at 6:18 PM.

Cryo techs at CHL reported, at 7:41 PM, seeing vibration spikes on their turbines at the same time as the quench. The Duty Electrician had turned off the Casey Pond water pumps for maintenance and that caused a rise in oil temperature. TeV experts did not believe this had anything to do with the quench.

TeV experts began a study period at 8 PM.

At 9:12 PM, a TeV vertical separator sparked. It sparked again ten minutes later. Operators talked with a TeV expert.

Operations loaded a 36x0 study-store into the TeV at 10:55PM.

At 11:57 PM, the TeV suffered a sector A1 quench. Operators discovered a separator that had gone to a wrong voltage and had tried to abort the beam.

Sunday August 17
The midnight shift began with the TeV recovering form a quench, with Pbar stacking, and with MiniBooNE taking beam.

A Linac RF station (LRF1) tripped off at 12:42 AM and then again fifteen minutes later. It reset okay both times.

Operations loaded at 4x0 study-store at 2:40 AM.

Operation put the TeV into shot setup at 3:11 AM and then established store 2916 at 5:22 AM. Its initial luminosity was 3.6E31. Pbar resumed stacking.

Booster RF stations 5&6 tripped off at 6:32 AM due to a high vacuum pressure alarm. Operations reset the pumps and turned the stations back on.

At 11:05 AM, Booster tripped off numerous times due to a bad monitoring device. Operators turned off the beam at 11:40 AM. With permission, operators replaced the device and beam resumed at 12:11 PM.

A Recycler expert began a study period at 3:16 PM.

A Switchyard expert began a study period at 7:46 PM.

At 11:28 PM, operators discovered a LCW trip holding off the power supplies for the magnets in the fixed target Meson area enclosure M05. The LCW pump was off and the water level was low. An expert thought there might be a leak in the system and came in to investigate.

NuMI studies began at 11:44 PM.

Monday August 18
The midnight shift began with operators searching for a LCW leak in the Meson area, with NuMI, Switchyard, and Recycler experts conducting studies, with Pbar stacking, and with MiniBooNE taking beam.

Operators discovered an LCW leak in the Meson MC6 enclosure. The expert investigated and found an unconnected magnet water hose. He valved off the hose.

At 1:42 AM, MiniBooNE began tripping off. Later, operators discovered that the NuMI studiers were doing things to the beam that was causing this problem.

The Meson LCW expert had the system back up and running by 3:33 AM.

At 3:38 AM, the Collider Detector Facility (CDF) suffered a device trip due to colliding beam problems. Later, operators discovered a CAMAC crate in service building E3 with no power. Operators replaced the power supply at 5:18 AM.

Operations terminated store 2916 at 5:02 AM and then put the TeV into shot setup at 7:46 AM.

Today's Plan
The Run Coordinator reported that we had an integrated luminosity of approximately 6 pb-1 for last week, which brings us close to our goal for the fiscal year of integrating 225 pb-1. The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/18/2003   email Fermilab